What is the wavelength of the x-ray emitted from a Tungsten target (Z=74) when an electron moves from the M-shell (n=3) to the K-shell (n=1)
1
Expert's answer
2020-06-08T10:17:35-0400

The energy of the electron when it moves from shell "n_1=3" to "n_2=1":


"E=(2.179\cdot10^{-18}\text{ J})Z^2\bigg(\frac{1}{n_2^2}-\frac{1}{n_1^2}\bigg)."

The wavelength of the x-ray emitted from a Tungsten target is


"\lambda=\frac{hc}{E}=\frac{hc}{(2.179\cdot10^{-18}\text{ J})Z^2\bigg(\frac{1}{n_2^2}-\frac{1}{n_1^2}\bigg)}=\\\space\\
=444\cdot10^{-6}\text{ m}."